XYZ Co. has estimated the cost of customer coordination activity as $1,000 per batch and the cost of inspection activity as $500 per inspection. The annual sales of 100,000 units of Product P is currently being produced in 10 batches of 10,000 units each and each batch requires 2 inspections, one at the start of production of the batch and the other at the end of production of the batch. The only customer for Product P has agreed to next year’s requirement being manufactured in 4 batches of 25,000 units each. Assume that all the costs in the two activity pools are controllable. The annual cost savings that management should aim to achieve at the coordination activity and inspection activity pools based on the customer’s new requirement plan is:

$8,000

$0

$20,000

$12,000

None of the above

$8,400
Answer #1

Calculate cost

Old New
coordination activity 1000*10 = 10000 1000*4 = 4000
Inspection activity 500*20 = 10000 500*8 = 4000
Total 20000 8000

So answer is d) $12000